# A PAC-Bayesian Framework for Optimal Control with Stability Guarantees

### v1

@@ I Introduction
Stochastic Nonlinear Optimal Control (SNOC) involves minimizing a cost function that averages out stochastic uncertainties affecting the nonlinear system dynamics, such as process noise and disturbances. Beyond the Linear Quadratic Gaussian scenario, this problem is often intractable and is typically replaced by minimizing the average cost across a dataset of sampled uncertainties, namely, the empirical cost. However, this approximation method leads to the challenge of determining generalization bounds, i.e., to upper-bound the true (uncomputable) cost using its (computable) empirical version. A remarkably similar problem has been thoroughly explored in Machine Learning within the scope of Probably Approximately Correct (PAC)-Bayesian learning [1]. Here, the generalization error of predictors drawn from an arbitrary posterior distribution is upper-bounded with a high probability, either in expectation or for a single draw. This upper bound comprises the empirical cost and a discrepancy measure between the posterior and a prior distribution. However, PAC-Bayesian bounds have never been exploited for the analysis of SNOC problems, with the exception of [2], where it was employed in a different context, focusing on zero-shot generalization to new environments. Related contributions utilize PAC-Bayes for learning control policies in Markov Decision Processes [3, 4]. However, these approaches assume finite action or state spaces, which might not be well-suited for SNOC. One reason for the limited application of classic PAC-Bayesian bounds in control is that they require the use of aggregated predictors, resulting from weighted averaging of all predictors according to the posterior distribution. In the SNOC context, such aggregation would require integrating actions from all possible controllers according to the posterior, a computationally challenging and impractical approach. Instead, we adopt a less commonly employed variant of PAC-Bayes bounds that apply to randomized predictors [1], which are single predictors drawn from the posterior. This allows us to establish a generalization bound when deploying a single controller sampled from the posterior, which is more practical. The proposed PAC-Bayesian bound provides an alternative to existing methods for validating certificates of learned controllers, such as those based on mixed-integer programs or satisfiability-modulo-theory certificates [5, 6].
@@
Recently, there has been a significant advancement on using neural networks in control loops [7, 8, 9, 10, 11, 12] or for providing stability certificates [13, 14]. In particular,[7, 8, 9] allow using broad families of nonlinear controllers (e.g., based on neural networks) for guaranteeing closed-loop stability while enabling the optimization of very general cost functions. The flexibility in selecting the cost function allows for promoting a broad spectrum of closed-loop specifications, ranging from safety to minimizing the L2-gain, among others. In this work, we exploit [7], which proposes an unconstrained parametrization of all and only stabilizing controllers for a given system. The controller implementation leverages Deep Neural Networks (DNNs) from [15] which are flexible architectures capable of achieving complex tasks. Since [7] guarantees closed-loop stability for any parameter choice, every posterior distribution over the induced space of parameters will exclusively assign non-zero probability to stabilizing controllers.
@@ I-A Contributions
We develop a novel framework for establishing a generalization bound for SNOC cost based on randomized bounds in PAC-Bayes theory, a class of bounds not previously exploited in control design. Based on the derived bound, we provide a new control policy design algorithm, which involves sampling the controller from an inferred posterior distribution. We leverage [7] to define this posterior distribution only over stabilizing controllers for a given system, enabling the deployment of our probabilistic algorithm while ensuring stability by design. To our knowledge, this is the first PAC-Bayesian control design algorithm with stability guarantees. Finally, we employ Stein Variational Gradient Descent (SVGD) [16] to make sampling the controller from the posterior numerically tractable. We demonstrate the feasibility of our method through two examples. First, we showcase how our approach can incorporate prior coarse knowledge into control design using a toy example. Then, we demonstrate the potential of our approach in a cooperative robotic task.
@@ III Generalization bounds
In this section, we establish a generalization bound on the true cost and use it for proposing a control design algorithm.
@@ III-A PAC-Bayes generalization bound
Assumption 1 may not apply to a wide range of SNOC problems. For instance, even a simple quadratic stage cost can lead to an unbounded $L$ if the closed-loop system is unstable. Therefore, we propose to map any unbounded cost function to $[0,C)$ by applying the following transformation:
@@
where $\tanh$ represents the hyperbolic tangent function and $\gamma\in\mathbb{R}^{+}$ is a suitable constant. The transformed cost grows almost linearly with the original cost between 00 and $\gamma$, and then smoothly saturates at the maximum value of $C$. In general, we expect that the stage cost, $l$, decreases over time. Therefore, the cost at the initial nominal state, $l(\bar{x},0_{m})$, is a reasonable selection for $\gamma$. This choice retains the transformed cost predominantly within the linear region rather than in the saturation region. In the sequel, we apply (7) to any unbounded FH cost (5) used in the SNOC problem (3), whenever necessary.
@@ III-B Optimal posterior
The bound (12) is more valuable when its right-hand side is smaller. This can be achieved by tuning $\lambda>0$, which according to Theorem 1, can be freely selected. However, computing the term $-1/\lambda\ln Z_{\lambda}(\mathcal{P},\mathbb{S})$ in (12) can be challenging due to its complexity. Under Assumption 1, this term is in $[0,C]$. Therefore, we replace $-1/\lambda\ln Z_{\lambda}(\mathcal{P},\mathbb{S})$ with $C$ to obtain an upper bound on (12). Below, we propose selecting $\lambda$ based on this looser upper bound.
@@ IV Learning stabilizing controllers
In order to guarantee closed-loop stability, we leverage the control design approach in [7]. In general, stability can be imposed in an SNOC problem (3) in two ways. The first one is to resort to constrained optimization approaches that ensure global or local stability by enforcing appropriate Lyapunov-like inequalities during optimization [14, 10, 11]. However, enforcing such constraints often becomes a computational bottleneck in complex applications. This issue has been highlighted in [5], discussing that popular stability verification methods are feasible for controllers with an order of $10^{2}$ parameters. The second possibility is to use unconstrained optimization approaches searching within classes of control policies with built-in stability guarantees [8, 9, 7]. These methods allow learning controllers through optimization algorithms based on gradient descent, without imposing further constraints to ensure stability. Additionally, unconstrained approaches usually have a much lower computational burden than constrained ones. In Section V-B, we show that unconstrained approaches also enable the use of algorithms for sampling the posterior, such as SVGD. Moreover, in Section VI-B, we demonstrate the scalability of these approaches by training a stabilizing controller with more than $10^{4}$ parameters.
In this work, we follow [7], which proposes an unconstrained parametrization of all and only stabilizing nonlinear control policies for a nonlinear time-varying system, as per (1), exploiting the internal model control framework [18]. This method is based on choosing a unique free operator, denoted as $\mathcal{M}_{\infty:0}$. The main result is that, assuming $\ell_{p}$-stability of system (1), the closed-loop map is $\ell_{p}$-stable if and only if the operator $\mathcal{M}_{\infty:0}$ is also $\ell_{p}$-stable. [^1] For ease of reading, we neglect the base controller in [7]; assuming that, if needed, it is already included in the system dynamics $f_{t}$. The control action $u_{t}$, for any $t\in\mathbb{N}_{0}$, can be obtained as:
@@
Implementing (14) requires parameterizing $\mathcal{M}_{\infty:0}$. This is addressed in [7] by adopting the Recurrent Equilibrium Networks (RENs) of [15], which allows representing a broad set of $\ell_{2}$-stable operators. Below, we briefly summarize this approach and refer the reader to [15] for a more comprehensive explanation. The resulting parametrized operator, :$\mathcal{M}^{\theta}_{\infty:0}$, is a dynamical system, given by:
@@ VII Conclusions and future work
This paper presents a novel generalization bound for the SNOC cost, drawing upon randomized bounds within the PAC-Bayes theory. By combining the derived bound with the unconstrained parametrization of stabilizing controllers proposed in [7], we propose an innovative control policy design algorithm suitable for addressing SNOC problems.
Possible directions for future work are: exploring model selection techniques among a set of sampled controllers to enhance performance while maintaining valid generalization bounds, and utilizing the knowledge obtained from existing controllers for a given system to set the prior distribution.

### v3

@@ I Introduction
Stochastic Nonlinear Optimal Control (SNOC) involves minimizing a cost function that averages out stochastic uncertainties affecting the nonlinear system, such as process noise and disturbances. Beyond the linear quadratic Gaussian scenario, this problem is often intractable and is typically replaced by minimizing the average cost across a dataset of sampled uncertainties, namely, the empirical cost. However, this approximation leads to the challenge of determining generalization bounds, i.e., to upper-bound the true (uncomputable) cost using its (computable) empirical version. A remarkably similar problem has been thoroughly explored in Machine Learning within the scope of Probably Approximately Correct (PAC)-Bayesian learning [1]. Here, the generalization error of predictors drawn from an arbitrary posterior distribution is upper-bounded with a high probability, either in expectation or for a single draw. This upper bound comprises the empirical cost and a discrepancy measure between the posterior and a prior distribution.
PAC-Bayesian bounds have never been exploited for the analysis of SNOC problems, with the exception of [2], where it was employed in a different context, focusing on zero-shot generalization to new environments. Related contributions utilize PAC-Bayes for learning control policies in Markov Decision Processes [3, 4]. However, these approaches assume discrete action or state spaces, which might not be well-suited for SNOC. One reason for the limited application of classic PAC-Bayesian bounds in control is that they require the use of aggregated predictors, resulting from weighted averaging of all predictors according to the posterior distribution. In the SNOC context, such aggregation would require integrating actions from all possible controllers according to the posterior, a computationally challenging and impractical approach. Instead, we adopt a less commonly employed variant of PAC-Bayes bounds that apply to randomized predictors [1], which are single predictors drawn from the posterior. This allows us to establish a generalization bound when deploying a single controller sampled from the posterior, which is more practical. The proposed PAC-Bayesian bound provides an alternative to existing methods for validating certificates of learned controllers, such as those based on mixed-integer programs or satisfiability-modulo-theory certificates [5, 6].
@@
Recently, there has been a significant advancement in using neural networks in control loops [7, 8, 9, 10, 11, 12, 13] or for providing stability certificates [14, 15]. In particular,[7, 8, 9, 10] allow using broad families of nonlinear controllers based on neural networks to guarantee closed-loop stability while optimizing general cost functions. The flexibility in selecting the cost function enables promoting a broad spectrum of closed-loop specifications, ranging from safety to minimizing the L2-gain, among others. In this work, we exploit the Performance-Boosting Control architecture, an unconstrained parametrization of all and only stabilizing controllers for a given system [7, 8]. Their implementation leverages Deep Neural Networks (DNNs) from [16] which are flexible models capable of achieving complex tasks. Since [7, 8] guarantee closed-loop stability for any parameter choice, every posterior distribution over the induced space of parameters will exclusively assign non-zero probability to stabilizing controllers.
@@ I-A Contributions
We develop a novel framework for establishing a generalization bound for SNOC cost based on randomized bounds in PAC-Bayes theory, a class of bounds not previously exploited in control design. Based on the derived bound, we provide a new control policy design algorithm, which involves sampling the controller from an inferred posterior distribution. We define this posterior distribution only over stabilizing controllers for a given system, enabling the deployment of our probabilistic algorithm while ensuring stability by design. To our knowledge, this is the first PAC-Bayesian control design algorithm with stability guarantees. Finally, we employ Stein Variational Gradient Descent (SVGD) [17] to make sampling the controller from the posterior numerically tractable. We demonstrate the feasibility of our method through two examples. First, we showcase how our approach can incorporate prior coarse knowledge into control design using a toy example. Then, we demonstrate the potential of our approach in a cooperative robotic task.
@@ III Generalization bounds
In this section, we establish a generalization bound on the true cost and use it to propose a control design algorithm.
@@ III-A PAC-Bayes generalization bound
Assumption 1 may not apply to a wide range of SNOC problems. For instance, even a simple quadratic stage cost can lead to an unbounded $L$ if the closed-loop system is unstable. We propose to map any unbounded cost function to $[0,C)$. To achieve this, we choose the transformation below:
@@
where $\tanh$ represents the hyperbolic tangent function and $\gamma\in\mathbb{R}^{+}$ is a suitable constant. This transformed cost grows almost linearly with the original cost between 00 and $\gamma$, and then smoothly saturates at the maximum value of $C$. In general, we expect that the stage cost, $l$, decreases over time. Therefore, the cost at the initial nominal state, $l(\bar{x},0_{m})$, is a reasonable selection for $\gamma$. This choice retains the transformed cost predominantly within the linear region rather than in the saturation region. In the sequel, we apply (7) to any unbounded FH cost (5) used in the SNOC problem (3), whenever necessary.
@@ III-B Optimal posterior
The bound (12) is more valuable when its right-hand side is smaller. This can be achieved by tuning $\lambda>0$, which according to Theorem 1, can be freely selected. However, computing the term $-1/\lambda\ln Z_{\lambda}(\mathcal{P},\mathbb{S})$ in (12) is often challenging as it involves calculating an expectation, as per (11). Under Assumption 1, this term is in $[0,C)$. Therefore, we replace $-1/\lambda\ln Z_{\lambda}(\mathcal{P},\mathbb{S})$ with $C$ to obtain an upper bound on (12). Below, we propose selecting $\lambda$ based on this looser upper bound.
@@ IV Learning stabilizing controllers
In order to guarantee closed-loop stability, we leverage the Performance-Boosting Control design approach of [7, 8]. In general, stability can be imposed in an SNOC problem (3) in two ways. The first one is to resort to constrained optimization approaches that ensure global or local stability by enforcing appropriate Lyapunov-like inequalities during optimization [15, 11, 12]. However, enforcing such constraints often becomes a computational bottleneck in complex applications. This issue has been highlighted in [5], discussing that popular stability verification methods are feasible for controllers with an order of $10^{2}$ parameters. The second possibility is to use unconstrained optimization approaches searching within classes of control policies with built-in stability guarantees [9, 10, 7, 8]. These methods allow learning controllers through optimization algorithms based on gradient descent, without imposing further constraints to ensure stability. Additionally, unconstrained approaches usually have a much lower computational burden than constrained ones [7, 8]. In Section V-B, we show that unconstrained approaches also enable the use of algorithms for sampling the posterior, such as SVGD. Moreover, in Section VI-B, we demonstrate the scalability of these approaches by training a stabilizing controller with more than $10^{4}$ parameters.
In this work, we follow [7, 8], which propose an unconstrained parametrization of all and only stabilizing nonlinear control policies for a nonlinear time-varying system, as per (1), exploiting the internal model control framework [19]. This method is based on choosing a unique free operator, denoted as $\mathcal{M}_{\infty:0}$. The main result is that, assuming $\ell_{p}$-stability of system (1), the closed-loop map is $\ell_{p}$-stable if and only if the operator $\mathcal{M}_{\infty:0}$ is also $\ell_{p}$-stable. [^1] For ease of reading, we neglect the base controller in [7]; assuming that, if needed, it is already included in the system dynamics $f_{t}$. The control action $u_{t}$, for any $t\in\mathbb{N}_{0}$, can be obtained as:
@@
Implementing (14) requires parameterizing $\mathcal{M}_{\infty:0}$. This is addressed in [7, 8] by adopting the Recurrent Equilibrium Networks (RENs) of [16], which allows representing a broad set of $\ell_{2}$-stable operators. Below, we briefly summarize this approach and refer the reader to [16] for a more comprehensive explanation. The resulting parametrized operator, :$\mathcal{M}^{\theta}_{\infty:0}$, is a dynamical system, given by:
@@ VII Conclusions and future work
This paper presents a novel generalization bound for the SNOC cost, drawing upon randomized bounds within the PAC-Bayes theory. By combining the derived bound with the unconstrained parametrization of stabilizing controllers proposed in [7, 8], we propose an innovative control policy design algorithm suitable for addressing SNOC problems.
Possible future directions are: further analyzing the bound’s asymptotic behavior, exploring model selection techniques among a set of sampled controllers while maintaining valid generalization bounds, and utilizing the knowledge obtained from existing controllers to set the prior distribution.